OpenStudy (anonymous):

if f(x)=2x, g(x)= 2x-1, and h(x)= square root of x, find (fogoh)(9)

OpenStudy (anonymous):

\[h(x)=\sqrt{x}\] \[h(9)=\sqrt{9}= 3\]

OpenStudy (anonymous):

Now you replace x=3 in g(x). \[g(x)=2x-1\] \[g(3)= 2(3)-1= 5\]

OpenStudy (anonymous):

Finally, replace x=5 in f(x) \[f(x)=2x\] \[f(5)= 2(5)= 10\]

OpenStudy (anonymous):

oh okay! i see... so do i then multiply them together? or add them?

OpenStudy (anonymous):

No. 10 is the answer. Actually, did them separately for you. Otherwise, it's like this: \[f(g(h(x)))= 2[2\sqrt{9}-1]\]

OpenStudy (anonymous):

But you get the same answer i.e. 10.

OpenStudy (anonymous):

oh okay!! thank you so much!!!!

OpenStudy (anonymous):

yw! :)